Please help, I thought it was true because the Neutron to Proton ratio is 1. Maybe I'm missing something. Please help and thank you.

162Tl is expected to be stable. Selected Answer: Correct Answer: B. True A. False
Answer #1

For atomic number less than 20, n/p ration is 1
For greater atomic number, n/p ratio should be greater than 1

Atomic number of Tl is 81
So, p = 81

Number of neutron, n = mass number - atomic number
= 162 - 81
= 81

n/p = 81/81 = 1:1

Since atomic number is greater than 1, this is not stable

Answer: false